OpenStudy (anonymous):

An object at a position s(0) = 4 begins to move at a velocity v(t) = e^(t/5). Find its position at t = 10.

OpenStudy (anonymous):

okay list all important information

OpenStudy (anonymous):

s(0) = 4 what?

OpenStudy (anonymous):

that's a strange velocity laughing out loud

OpenStudy (anonymous):

\[s(t) = \int\limits v(t)dt\]

OpenStudy (anonymous):

\[= 5e^{\frac{ t }{ 5 }} + c\]

OpenStudy (anonymous):

use that 4 for your c

OpenStudy (anonymous):

then solve for s(10)

OpenStudy (anonymous):

I hope @sourwing approves of my solution

OpenStudy (anonymous):

it's a strange unitless physical application

OpenStudy (anonymous):

these are the answer choices: e2 - 1 e2 - 5 5e2 - 1 5e2 - 5

OpenStudy (abb0t):

0<t<10

OpenStudy (abb0t):

is what you should be integrating.

OpenStudy (anonymous):

what about c, then?

OpenStudy (abb0t):

There is no constant, you're integrating from \(\sf \color{red}{ 0<t<10}\)

OpenStudy (anonymous):

so the s(0)=4 was unnecessary info?

OpenStudy (abb0t):

i didn't even see that. Uh, so you have \(\sf \color{red}{4=5+C}\), solve for \(\sf \color{blue}{C}\)

OpenStudy (abb0t):

But your answer choices are wrong, if that is the case.

OpenStudy (abb0t):

It would not be 5e\(\sf ^2-1\)

OpenStudy (abb0t):

Where is the e\(^2\) coming from?!?!?!

OpenStudy (anonymous):

so the answer is 5e^2−1. because you solve the integral, then find c which is -1.

OpenStudy (anonymous):

where is 5 coming from when solving C?

OpenStudy (anonymous):

that was incorrect. i was solving the definite integral someone gave and plugged that it. ignore that whole 5e^2−5 thing whoops.

OpenStudy (anonymous):

c=4 is not an unnecessary information s(0) means that is your initial position, that is a t=0 you are at 4

OpenStudy (anonymous):

the derivative of position function is velocity and the anti-derivative (integral) of velocity is position function we took the integral of velocity and in place of C we put the initial position

OpenStudy (anonymous):

wait how is c=4? 5e^t/5 is the anti derivative of the given equation, right? if you plug t=0 into that anti derivative, wouldn't you get 5? e^0=1. therefore c=-1?

OpenStudy (anonymous):

\[s(t)=\int\limits v(t)dt= \int\limits e^{\frac{ t }{ 5 }}dt=5e^{\frac{ t }{ 5 }}+C\] to solve for C \[s(0)=5e^{\frac{ t }{ 5 }}+C=4\] looking at the exponent \[\frac{ 0 }{ 5} = 0 \rightarrow 5e^0 = 5*1=5\] 5+C = 4 C = -1 evaluate at s(10) meaning plug 10 in everywhere you see a t

OpenStudy (anonymous):

so the answer is 5e^2−1

OpenStudy (anonymous):

\[s(t)=5e^{\frac{ t }{ 5 }}-1\]

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

the answer has been given you just need to apply yourself a little bit